OpenStudy (anonymous):

Find the smaller angle between vectors a and b a= <3,2,1> b=<-3,0,-1>

OpenStudy (anonymous):

so this is what i have so far, and i'm not sure if i did it correctly or not... aob = ||a|| ||b|| CosTheta -10 = \[-10= \sqrt(14)\sqrt(10) Cos \theta\]

OpenStudy (anonymous):

so i get \[\cos \theta = -10/ 11.8\]

OpenStudy (anonymous):

if i arctan this i get 32.4, but apparently that's not the right answer? o.O

OpenStudy (anonymous):

i mean, arc cos lol

OpenStudy (anonymous):

yes.you are right.

OpenStudy (anonymous):

my teacher marked it wrong :( i think i may have done something wrong somewhere? o.O

OpenStudy (cwrw238):

i got 32.06 degrees

OpenStudy (anonymous):

i think you may have done some calc. mistakes

OpenStudy (anonymous):

wait, i got 32.06 on my calculator lol but marked down 32.4 would she seriously mark me off 3 points because i'm off .34 points..???

OpenStudy (cwrw238):

wait isnt aob = -8 ? no - forget that -10 is right

OpenStudy (cwrw238):

yea 32.06 is correct

OpenStudy (cwrw238):

3 points off is a bit harsh - marks are usually awarded for method and accuracy and you had the method right

OpenStudy (anonymous):

is it the smaller of the two angles? maybe that's why it's wrong? is it possible to have a smaller angle?

OpenStudy (anonymous):

no.the other one is